OpenStudy (anonymous):

system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

FAN AND MEDAL?!

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

lol hey yea sry i was on another window

OpenStudy (anonymous):

oh its okay! can you help me?

OpenStudy (anonymous):

yep, for the sum of the two equations, just add the left side of the first equation with that of the second, same thing with the right side

OpenStudy (anonymous):

well i already solved the first equation and got (1,-1). was i wrong to solve it?

OpenStudy (anonymous):

you want me to set them both equal to eachother?

OpenStudy (anonymous):

umm no not yet.

OpenStudy (anonymous):

okay so what do i do

OpenStudy (anonymous):

how did u solve the first equation? when u plug in (1,-1), it doesn't work

OpenStudy (anonymous):

i multiplied the first equation (8x+2y=9) by negative one so i could cancel out the Y's. then i just solved using elimination for x then got an x value of 1 and plugged it back into the first equation to solve for y

OpenStudy (anonymous):

the Y's wont cancel out, when u multiple +2y by -1, u get -2y, this won't cancel out the +5y from the second equation

OpenStudy (anonymous):

oh the first equation is 8x + 5y = 9 not 8x+2y=9

OpenStudy (anonymous):

oh wow... i copied down a 2. my bad .!!

OpenStudy (anonymous):

so do you want me to solve the two equations first?

OpenStudy (anonymous):

its ok, umm...yea, tell me wut u get

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so now that there isnt 2s how can i solve them?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

u can multiply the first equation by 2, and the second equation by 5, this is to get the LCD (least common denominator), ull then get +10y for both equation. now all u hafta do is multiple one of the equation by -1, and the -10y will cancel out the 10y, u can then solve the now-isolated x

OpenStudy (anonymous):

u can do the same LCD method for x, and isolate then solve for y

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

for x i got -2

OpenStudy (anonymous):

for y i got 5

OpenStudy (anonymous):

yes there we go, correct!

OpenStudy (anonymous):

okay cool! so now what do we do to find the equivalent equations? thats the most confusing part

OpenStudy (anonymous):

hmm what do you mean equivalent equations? like how 2y = 3 is the same thing as 4y = 6?

OpenStudy (anonymous):

yeah. like back up at the problem it says Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yea im here so sry man :P monitoring the chat...

OpenStudy (anonymous):

its okay haha. so what do i do

OpenStudy (anonymous):

okay, so they want us to combine the first equation with a multiple of the second equation

OpenStudy (anonymous):

how dod i do that? mutliply the second equation by 2 and add it to the first one? wouldnt that only give us one equation?

OpenStudy (anonymous):

well, u replace the first equation, the second equation u keep

OpenStudy (anonymous):

okay so what dod i multiply the second equation by?

OpenStudy (anonymous):

i guess any constant should do

OpenStudy (anonymous):

so 2?

OpenStudy (anonymous):

lol but i dont get the point thou...this question is weirdly phrased

OpenStudy (anonymous):

and i would have 6x+4y=8

OpenStudy (anonymous):

and i just add that to the first equation?

OpenStudy (anonymous):

lol i guess so yes

OpenStudy (anonymous):

so the new first equation i got is 14x+9y=17

OpenStudy (anonymous):

u multiplied the second equation by 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mmk, that's it?

OpenStudy (anonymous):

would my new equivalent equations be 14x+9y=17 and 3x+2y=4

OpenStudy (anonymous):

well after i multiplied the second equation by 2 i added it to the first equation

OpenStudy (anonymous):

hmm is this homework or like a quiz/test?

OpenStudy (anonymous):

its a study guide for a test

OpenStudy (anonymous):

but i need the right answer

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

okay so ???

OpenStudy (anonymous):

is that new set of equivalent equations right?

OpenStudy (anonymous):

OHHH okay so according to Mathwords, an equivalent system of equations is basically another set of equations with the same solution (a point), which, in this case, is (-2, 5)

OpenStudy (anonymous):

okay so is my new system i just got not correct?

OpenStudy (anonymous):

testing it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yyes it works, oh wow this is such an inconvenient way to do this lol never seen anything like this

OpenStudy (anonymous):

i KNOW! very complicated! thank you so much. can you help me do another?

OpenStudy (anonymous):

yea of course

OpenStudy (anonymous):

okay hold on im going to attach the file so itll be easier to see

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

/Users/lyndsay/Desktop/Screen Shot 2014-05-23 at 11.54.37 PM.png

OpenStudy (anonymous):

ugh sorry trying to figure out how to attch it

OpenStudy (anonymous):

LOL yea

OpenStudy (anonymous):

Can you open it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

well for part a, ur basically just substituting the two original functions, this is an example of an equivalent system of equations, and as we're already talked about, they will have the same solution(s)

OpenStudy (anonymous):

so i substitute what? The question itself is worded weird

OpenStudy (anonymous):

ur substiuting the 2 y's in the second equation

OpenStudy (anonymous):

y = y, replaced y's, u get the second equation, get it? sry i cant explain this smoothly lol

OpenStudy (anonymous):

so for part A the answer would be because they are both equivalent equations?

OpenStudy (anonymous):

yea, don't forget to mention the word "sustitute"

OpenStudy (anonymous):

so i should write They are both equivalent equations because they are substituted into the equation? or how should i word it?

OpenStudy (anonymous):

They have the same solutions because they are an example of an equivalent equation system, having been substituted by the previous 2 y equations

OpenStudy (anonymous):

okay so now for part b?

OpenStudy (anonymous):

u know how 4 is 2 ^ 2? and so 4^x can be rewritten as 2 ^2x, do u understand this rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so u also know that if 2 exponential functions have the same root, their powers are equal, yes?

OpenStudy (anonymous):

yes then you can set the exponents equal and solve for x right?

OpenStudy (anonymous):

yes, but remember to turn the base 4 into base 2

OpenStudy (anonymous):

okay. so how do i do the table of values?

OpenStudy (anonymous):

or wait what does it mean by make tables??

OpenStudy (anonymous):

yes, so just plug in x values that are within the given range

OpenStudy (anonymous):

okay so the equation would be 2^2x=2^x+3??

OpenStudy (anonymous):

then i just make a table to plug in values?

OpenStudy (anonymous):

yes, so set 2x equal to x + 3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

lol wth...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i set 2x=x+3 and solved for x

OpenStudy (anonymous):

lol yea i know and ur correct, just the table part...

OpenStudy (anonymous):

im confused on how to plug in the values

OpenStudy (anonymous):

like idk how i should make the table nd what to plug in where

OpenStudy (anonymous):

oh maybe u weren't suppose to solve it via base-changing, but plug and chug..plug in x for values between -3 and 3, and use the coordinates to make the table. when the 2 equation equals, there's ur solution (although we found it already lol derp)